Calculate the pressure change if two moles of
steam at 5 x 102K occupying 0.03 m3of volume
is heated up to 1 x 103 K at constant volume.
Assume that steam behaves as Van der Waals
gas. It is given that
1
Expert's answer
2019-06-03T05:48:18-0400

Solution.

V = const.


"\frac{p1}{T1} = \frac{p2}{T2}"

"p \times V = n \times R \times T"

"p1 = \frac{n \times R \times T1}{V1}"

p1 = 277000 Pa

p2 = 554000 Pa


"\Delta p = p2-p1"

"\Delta p = 277000 Pa"

Answer:

"\Delta p = 277000Pa"